OpenStudy (anonymous):

the amount of money in an acount with countinously compounded interest is given by the formula A=pe^rt , where p is the principle, r is the anual interest rate, and t is the time in years. calculate to the nearest tenth of a year how long it takes for an acount of money to double if interest is compounded continusly at 7.5%

OpenStudy (anonymous):

You want the money in the account to double, which means \(A=2p\): \[2p=pe^{rt}\\ 2=e^{rt}\] The rate is 7.5%, or \(r=0.075\): \[\large2=e^{0.075t}\\ \large\ln 2=\ln\left(e^{0.075t}\right)\\ \large\ln 2=0.075t\ln e\\ \large\ln 2=0.075t\\ t=\cdots\]

OpenStudy (anonymous):

8.2 years?

OpenStudy (anonymous):

9.2 Years?

OpenStudy (wolf1728):

Or you could use this formula: Years = {log(total) -log(Principal)} ÷ log(1 + rate) where you would make total =2 and principal =1 Years = [log(2) - log(1)] /log (1.075) Years = [0.3010299957 -0] / 0.0314084643 Years = 0.3010299957 / 0.0314084643 Years = 9.5843589566 and I found the formula here: www.1728.org/compint2.htm

OpenStudy (anonymous):

it is definalty 9.2 years correct?

OpenStudy (wolf1728):

Natural log of 2 = 0.69314718056 0.69314718056 = .075*t t = 0.69314718056 / .075 t = 9.2419624075 (guess I made a mistake in my own calculations)

OpenStudy (anonymous):

A = Pe^(r)(t) 2 = e^0.075t ln2 = 0.075t 0.6931 = 0.075t t = 9.242 years

OpenStudy (wolf1728):

That seems right

OpenStudy (anonymous):

thank you wolfman!!! lol

OpenStudy (wolf1728):

thanks :-)

OpenStudy (wolf1728):

I found out what I did wrong. I did not use a continuously compounded interest rate and just used an annual rate. (I must read the questions more carefully).